help with this question please, clear steps. I will appreciate it!
If P is the parallelogram bounded by the lines x - y = 0, x + 2y = 0, x - y = 1 and + 2y 6 then: SI xy dA
Answer #1

p is the parologram bounded by the lines X-Y=ox+2y=0, x=y= 1 and x+2y= b We Arred to find 2324 62,2) 7-y=1 Sf xy da (C) Ssxy